Find the direction cosine (l, m, n) of vector r = 3i -2j + 6k b. Given that p = -12 and q = 1. Find a unit vector u such that │u│= 35 and u is the direction of p-3q 5 -1

1
Expert's answer
2022-06-15T14:47:54-0400

a.


"\vec{r} = 3\vec{i} -2\vec{j} + 6\vec{k}"

"|\vec{r}|=\sqrt{(3)^2+(-2)^2+(6)^2}=7"

"(l, m, n)=(\dfrac{3}{7}, -\dfrac{2}{7}, \dfrac{6}{7})"

b.


"\vec{u} =\langle k(-12-3(1)), k(5-1)\rangle"

"\vec{u} =\langle-15k, 4k\rangle"

"|\vec{u}|=\sqrt{(-15k)^2+(4k)^2}=35"

"241k^2=35^2, k>0"

"k=\dfrac{35}{\sqrt{241}}"

"\vec{u} =\langle-\dfrac{525}{\sqrt{241}}, \dfrac{140}{\sqrt{241}}\rangle"